OpenStudy (anonymous):

SOlve 13^4x = 7 for x. I got like a random negative decimal... what am i doing wrong?

OpenStudy (unklerhaukus):

\[13^{4x}=7\] is this your question/

OpenStudy (unklerhaukus):

or \(13^4x=7\)

OpenStudy (anonymous):

the first one...

OpenStudy (unklerhaukus):

what answer did you get/

OpenStudy (anonymous):

i got -3.241345587040701

OpenStudy (anonymous):

Take the natural log of both sides: 13^4x = 7 ln13^4x = ln7 by the properties of logarithms, we can pull down the exponent 4x: 4x(ln13) = ln7 4x = ln7/ln13 x = 4(ln7/ln13)

OpenStudy (anonymous):

the answer should be about 3.03....

OpenStudy (unklerhaukus):

your method is good @Hermeezey but in you last step you made a little mistake \[4x = \frac{\ln7}{\ln13}\] \[x=\frac{\ln7}{4\ln13}\approx3.03\]

OpenStudy (unklerhaukus):

i mean \(x\approx0.1897\)

OpenStudy (anonymous):

oh yea sorry, i get confused when i'm typing

OpenStudy (anonymous):

thanks :)

OpenStudy (unklerhaukus):

it is always good to check the solution by pluggin it into the original equation \[13^{4\times0.1897}\approx13^{3/4}\approx6.8\approx7\]